OpenStudy (anonymous):

Please help, I really need it: Compute the 8th derivative of...

OpenStudy (anonymous):

\[\large f(x)=\frac{\cos(3x^3)-1}{x^{4}}\]

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

a hint was to use the MacLaurin series for f(x)

OpenStudy (anonymous):

if you do the MacLaurin series for the cos(3x^3) you get:\[\Large \sum_{n=0}^{\infty} {\frac{(3x^{3})^{2n}}{(2n)!}}(-1)^{n}\]

OpenStudy (anonymous):

@ash2326 @estudier

OpenStudy (anonymous):

and then if you minus 1 (as per f(x)) and divide each term by x^4, then, when n=4 (the 8th derivative) i get:\[\Large \frac{\frac{3^{8}x^{24}}{8!}}{x^{4}}\]

OpenStudy (anonymous):

which would imply that\[f^{8}(0)=3^{8}\]

OpenStudy (anonymous):

(the question is to calculate the 8th derivative at x=0:\[f^{8}(0)=.....\])

OpenStudy (ash2326):

@remnant What's your f(x) in terms of maclaurin series of cos 3x^3?

OpenStudy (anonymous):

\[f(x) = \frac{\Large \sum_{n=0}^{\infty} {\frac{(3x^{3})^{2n}}{(2n)!}}(-1)^{n}-1}{x^4}\]

OpenStudy (ash2326):

Expand it and write the terms

OpenStudy (anonymous):

i did...and my fourth term (which has the 8th derivative in it) was:\[\Large \frac{\frac{3^{8}x^{24}}{8!}}{x^{4}}\]

OpenStudy (ash2326):

Nope, second term n=2 \[\frac{(3x^3)^4}{4!} \div x^4\] \[\frac{3^4 x^8}{4!}\] all terms on the left will get to 0 by 8th derivative, if you take 8th derivative of this term you'd get \[\frac{3^4 \times 8!}{4!}\] if you put x=0, only this will be left \[f_8(0)=\frac{3^4 \times 8!}{4!}\]

OpenStudy (anonymous):

surely according to my MacLaurin series for cos(3x^3), the eight derivative will happen when n=4 as cos skips one per one done

OpenStudy (ash2326):

@remnant 1) first step is to simplify f(x). You could write in series form, first two terms maybe 2) Next step is to notice what happens on derivative, start in steps of 1 first term is x^2 and second term is x^8, rest are higher powe terms obviously x^2 diminishes to 0 in 3rd derivative x^8 becomes 1 if you take 8 times derivative higher order terms will have x if you put x=0, only one term is left.

OpenStudy (ash2326):

I'll be back in 20 minutes, to explain you. Till then you try to check

OpenStudy (anonymous):

thanks, will do

OpenStudy (anonymous):

okay so i get the right answer now and I kinda get how/why i'm doing what i'm doing....when n=2 i get\[\frac{81x^{12}}{24x^{4}}\]and so then i get\[\frac{27}{8}x^8\] which if you take 8 derivatives of, you get\[(8!)\frac{27}{8} = 136080\]

OpenStudy (anonymous):

BUT....i have no idea why i would just randomly stop when n=2? ... is it cuz before then i would get 0 and after then i would still have some x's lying around

OpenStudy (amistre64):

when n=2, I get:\[\frac{81x^{12}-1}{24x^{4}}\]

OpenStudy (anonymous):

nah, the first term of the series is 1 so the -1 cancels that out

OpenStudy (amistre64):

ahhh :)

OpenStudy (amistre64):

i wish i had more time to focus on this, but trying to do some job related stuff.

OpenStudy (ash2326):

\[f(x)=\frac{\cos (3x^3)-1}{x^4}\] rewriting \(\cos (3x^3) \) in terms of its MacLaurin series \[\Large f(x)=\frac{(1-\frac {(3x^3)^2}{2!}+\frac {(3x^3)^4}{4!}-\frac {(3x^3)^6}{6!}....)-1}{x^4}\] \[\Large f(x)=\frac{(\cancel 1-\frac {(3x^3)^2}{2!}+\frac {(3x^3)^4}{4!}-\frac {(3x^3)^6}{6!}....)-\cancel 1}{x^4}\] \[\Large f(x)=\frac{(\frac {(3x^3)^2}{2!}+\frac {(3x^3)^4}{4!}-\frac {(3x^3)^6}{6!}....)}{x^4}\] \[\Large f(x)=\frac{\frac {3^2x^6}{2!}+\frac {3^4x^{12}}{4!}-\frac {3^6x^{18}}{6!}....}{x^4}\] \[\Large f(x)=\frac {3^2x^2}{2!}+\frac {3^4x^{8}}{4!}-\frac {3^6x^{14}}{6!}....\] Do you get this ?

OpenStudy (anonymous):

yip :) (thanks for writing that out :) )

OpenStudy (ash2326):

Now let's understand when a x term will become zero on differentiation Example: \[f(x)=x\] \[f'(x)=1\] \[f''(x)=0\] \[f(x)=x^2\] \[f'(x)=2x\] \[f''(x)=2\] \[f'''(x)=0\] so if the power of x is n, then on (n+1) th derivative the term will become zero Do you get this?

OpenStudy (anonymous):

yip :)

OpenStudy (ash2326):

\[\Large f(x)=\frac {3^2x^2}{2!}+\frac {3^4x^{8}}{4!}-\frac {3^6x^{14}}{6!}....\] here first term will become zero on 8th derivative second will become zero only on 9 th, higher power terms will have x term on 8th derivative \[f''''''''(x)=0+\frac{3^4 \times 8\times 7 \times 6\times 5\times 4 \times 3 \times 2\times 1}{4!}-\frac {3^6 \times 14\times 13\times ....7\times x^6}{6!}....\] now put x=0

OpenStudy (ash2326):

Only second term will be there, rest all will vanish \[f''''''''(x)=0+\frac{3^4 \times 8!}{4!}+0\]

OpenStudy (ash2326):

Sorry \[f''''''''(0)=0+\frac{3^4 \times 8!}{4!}+0\]

OpenStudy (anonymous):

aaaaaaah :)

OpenStudy (ash2326):

@remnant do you understand this?

OpenStudy (anonymous):

fully with you....just dont think i would ever "see" that

OpenStudy (ash2326):

you got it now?

OpenStudy (anonymous):

thanks for spending so much time going through the answer....was really really helpful :)

OpenStudy (anonymous):

yeah i got it :)

OpenStudy (ash2326):

Glad to help you:D